OpenStudy (gabylovesyou):

Given the functions j(x) = x2 - 9 and k(x) = -x + 7, which operation results in a 3rd degree polynomial?

OpenStudy (ybarrap):

Neither j(x) nor k(x) result in a 3rd degree; however, when you multiply them it will. How do we know that? Well, if you look at the highest degree of j(x), which is 2 and for k(x), which is 1, their product will result in a term involving the product of \(x^2\) in j(x) and \(x\) in k(x). And \(x^2\times x=x^3\), which is 3rd degree

OpenStudy (ybarrap):

This is what I'm saying, $$ (x^2-9)\times (-x+7)=-x^3+7 x^2+9 x-63 $$ Note the cubic term (i.e. 3rd degree) in the final result.

OpenStudy (ybarrap):

Implied here is your answer. What is the "operation" I used here?

OpenStudy (gabylovesyou):

so multiplication.... got it! thanks !

OpenStudy (mathstudent55):

Are there choices?

OpenStudy (ybarrap):

You're welcome